The worlds's strongest current, reaching 7 miles per hour, is the Saltstraumen in Norway. A motorboat can travel 54 miles downstream in the Saltstraumen in the same time it takes to go 18 miles upstream. What is the speed of the boat i still water?
Answer by jorel1380(3719)

Let n be the speed of the boat in still water; and the current is 7 mph. So:
54/n+7=18/n-7
18n+126=54n-378
36n=504
n=14 mph as the speed of the boat in still water. ☺☺☺☺